# A new approach to quark mass determination using the gradient flow

### v1

@@ 2.2 Expressing the VEVs in terms of scalar integrals
$S|_{\rm 2-loop}$ $=g_{0}^{2}N_{\mathrm{c}}C_{\mathrm{F}}\Big{[}8(d-2)I_{1}-32I_{2}-16I_{3}-16I_{4}$ (11)
$\qquad-8I_{5}+32I_{7}-4(d-2)I_{10}+16I_{11}\Big{]}\,,$
$R|_{\rm 2-loop}$ $=g_{0}^{2}N_{\mathrm{c}}C_{\mathrm{F}}\Big{[}16J_{1}+(4d-8)J_{2}+2(8-4d)J_{3}+(16d-32)J_{5}$
$\qquad-64J_{6}+32J_{7}-32J_{8}+16J_{9}-32J_{10}+64J_{11}+32J_{12}$
$\qquad-64J_{13}+16J_{14}+(8-4d)J_{15}-(8-4d)J_{16}-(8-4d)J_{17}$
$\qquad-(16-8d)J_{18}+(16-8d)J_{19}-32J_{20}\Big{]}\,.$
@@ 2.3.1 Laplace transform
where $v_{0}$ is a real number chosen such that the integral of eq. (14) converges for $v=v_{0}$. If $\hat{I}=\mathcal{O}(z^{a_{s}})$ for small $z$ and $I=\mathcal{O}(z^{-a_{l}})$ for large $z$, one may choose $-a_{l}<v_{0}<a_{s}$. This also implies that $\tilde{I}(v)$ develops singularities in the $v$-plane which encode the structure of the expansions of $I(z)$ in $z$ and in $1/z$, as we will now explain.
@@ 2.3.3 Two-loop examples
×$\quad{}\times\int_{0}^{1}\int_{0}^{1}\mathrm{d}u\,\mathrm{d}u^{\prime}\,u^{\epsilon}(4-u(1-u^{\prime})^{2})^{-\frac{3}{2}+v+\epsilon}{}_{2}F_{1}(1,v-\frac{1}{2};2-\epsilon,\frac{1}{4}u(1-u^{\prime})^{2}),$ (32)
@@
$v$ =±$=\pm\left(n-\frac{1}{2}\right)\qquad\text{and}\qquad v=n+\frac{3}{2}-\epsilon\,,\qquad n\in\mathbb{N}\,.$ (33)
@@
where $\mathcal{T}_{1,N}g_{a}(y)$ denotes the Taylor expansion of $g_{a}(y)$ around $y=1$ up to power $(1-y)^{N}$, $g_{a}^{(k)}(y)$ is the $k^{\text{th}}$ derivative of $g_{a}(y)$, and $N$ an integer part of $v-3/2$ (or $v-3/2\leq N<v-1/2$).
@@
with the integer $N=-v-3/2$. Note that, since the pre-factor only develops poles at positive $v$, the residues at negative $v$ are given entirely by the Taylor expansion coefficients $h^{(k)}(0)$ of the function $h(y)$ which can be easily evaluated analytically.
@@
This integral involves an additional complication, in that its Laplace transform contains singularities at $v=0$ and $v=-\epsilon$, meaning that $v_{0}$ in eq. (15) has to be chosen carefully. As argued in app. B, the choice $-\epsilon<v_{0}<0$ properly separates the small- and large-$(m^{2}t)$ expansions, assuming $\epsilon>0$.
@@ 2.4 Differential equations
$i(t)=\sum_{k=0}^{\infty}\left(\frac{a_{k}}{k+2\epsilon}m^{2k+3}t^{k+2\epsilon}+\frac{b_{k}}{k+\epsilon}m^{2(k-\epsilon)+3}t^{k+\epsilon}+\frac{c_{k}}{k}m^{2(k-2\epsilon)+3}t^{k}\right)+\text{const.}$. (45)
@@ 2.5 Numerical evaluation
Since we expect the asymptotic expansions of the integrals to fail in the intermediate region, we also evaluate the integrals numerically on a one-dimensional grid in $m^{2}t$. This is achieved with the help of the program ftint Harlander:2024vmn . It uses Schwinger parametrization of the momentum integrations in order to map the integrals onto a multi-dimensional hypercube. The integrand becomes the product of polynomials in the Schwinger and flow-time parameters, raised to $d$-dependent powers. Such integrals can be evaluated algorithmically by sector decomposition Binoth:2000ps ; Binoth:2003ak , which expresses them in terms of a Laurent series in $d-4$ whose coefficients are convergent integrals which can be evaluated numerically. For the sector decomposition and the numerical integration, ftint relies on pySecDec Borowka:2015mxa ; Borowka:2017idc ; Borowka:2018goh ; Heinrich:2023til . Evaluation of all integrals to a relative precision of $10^{-8}$ or better on a grid of 200 points takes only a few hours on a currently average computer.
@@ 2.6 Small-flow-time expansion
In addition to a direct calculation of the mass expansion outlined above, one can also determine the expansion order-by-order by employing the small-flow-time expansion on the operator level Luscher:2011bx as was briefly mentioned in ref. Lange:2021vqg . This effectively amounts to an operator product expansion with $t\to 0$. For $S(t)$ it reads
@@ 3.1 Perturbative results
The renormalized mass and coupling in the $\overline{\rm MS}$ scheme are given through
@@
$\overline{S}$ $:=\langle\bar{\chi}_{\mathrm{R}}(t,x)\chi_{\mathrm{R}}(t,x)\rangle\,,$ (57)
@@
Figure 1: The perturbative coefficients of the ratio of $\overline{S}$ (left) and those of $\overline{R}$ (right), obtained using two approaches, i.e., expansions and numerical evaluation, as a function of $\overline{m}^{2}t$, cf. eq. (59). The lines in the region $\overline{m}^{2}t<1$ correspond to the small-$\overline{m}^{2}t$ expansion (blue and purple lines) and those in $\overline{m}^{2}t>1$ to the large-$\overline{m}^{2}t$ expansion (red and orange lines). Since the $\mathcal{O}(\alpha_{\mathrm{s}})$ coefficient also depends on $\mu/\overline{m}$ in addition to $\overline{m}^{2}t$, we show results with a few choices: $\mu/\overline{m}=100$ (purple solid), $5$ (purple dashed), $1$ (purple dotted), and $\mu/\overline{m}=1$ (orange solid), $2$ (orange dashed), $1/2$ (orange dotted).
@@ 3.2 Lattice observables and uncertainty estimates for quark mass extractions
$\alpha_{\mathrm{s}}(M^{2}_{\mathrm{Z}})$ $=0.1180\,,$ (70)
$\overline{m}_{\mathrm{s}}(2~{}{\rm GeV})$ $=93.5\,\text{MeV}\,,$
$\overline{m}_{\mathrm{c}}(\overline{m}_{\mathrm{c}})$ $=1.2730\,\text{GeV}\,,$
$\overline{m}_{\mathrm{b}}(\overline{m}_{\mathrm{b}})$ $=4.183\,\text{GeV}\,.$
@@
Thus, we can determine $\overline{m}$ only up to ±$\pm\Delta\overline{m}$, where
@@
Figure 5: Estimate of the expected accuracy on mass determination at the current perturbative order $[\mathcal{O}(\alpha_{\mathrm{s}})]$. See the text for details.
In the case of the strange and charm quarks, where the small-$\overline{m}^{2}t$ is more relevant, we see the tendency that $\Delta\overline{m}/\overline{m}$ decreases towards smaller values of $t$, because the perturbative uncertainty decreases in this case. For the strange quark, the effect of the larger perturbative uncertainty for $r_{b}$ compared to $r_{c}$ is partly compensated by the larger derivative, leading to a similar expected accuracy on $\overline{m}$. In the case of the bottom quark, the figure on $r_{b}$ indicates the potential to extract its mass value at $1-2$ already at this perturbative order. We note that these are crude estimates, based on the scale variation around our central scale choices, eq. (68) or (69). Therefore they are understood as order-of-magnitude estimates.
Besides the perturbative uncertainty due to missing higher orders, we also have to take into account the uncertainties from the lattice determinations for a full uncertainty estimate. In addition, there could also be non-perturbative corrections which we discuss in the next section.
@@ 4 Discussion of non-perturbative effects
where $c_{0,1,2}$ are constants, one can see that the largest non-perturbative contribution, $c_{0}\Lambda_{\rm QCD}^{3}$, which is independent of $m$, disappears due to the derivative in $r_{c}$. Hence, the leading non-perturbative effect reduces to $\mathcal{O}(m\Lambda_{\rm QCD}^{2})$, whose magnitude relative to the perturbative effect $m/t$ reduces to $\mathcal{O}(t\Lambda_{\rm QCD}^{2})$, which is much smaller than one and can be well controlled. As for $1-R/(R|_{m=0})$, the leading non-perturbative effect is $\mathcal{O}(m\Lambda_{\rm QCD}^{3})$, which is smaller than that found in $R/(R|_{m=0})$, but still can be large compared with the perturbative effect $m^{2}/t$. Based on these observations, we conclude that $r_{c}$ is the most suitable quantity for light quark mass determination among $r_{a,b,c}$ and $1-r_{b}$.
@@ 5 Conclusions
We proposed a method to determine quark masses using the gradient flow by considering ratios of VEVs of flowed quark-bilinear operators. They are scheme and gauge invariant and can be calculated both on the lattice and in perturbation theory. We evaluated these VEVs perturbatively up to two loops, i.e. $\mathcal{O}(\alpha_{\mathrm{s}})$, including finite mass effects. Compared to the already known perturbative results in the small-$\overline{m}^{2}t$ limit, we found that the higher-order terms of the small-$\overline{m}^{2}t$ expansion, the large-$\overline{m}^{2}t$ expansion, and full mass dependence are crucial for studying particularly charm and bottom quarks. The perturbative behavior of the VEV ratios exhibit good perturbative behavior. We also estimated the expected accuracy achieved in the mass determination from the perturbative viewpoint.
@@ Appendix B Choosing $v_{0}$ for the integral $J_{17}$
In the first inequality, we used $1/[(p-k)^{2}+m^{2}]<1/(p-k)^{2}$. The bound (83) is valid as long as $42<d<$ 4; the $k$-integral is UV divergent for $d\geq 4$ and infrared divergent for $d\leq 2$. Therefore, we assume this range of $d$ and thus $\epsilon>0$. $C$ and $C^{\prime}$ are positive constants. From this bound, we can see that the $v=-\epsilon$ singularity, which gives $\mathcal{O}(m^{-2\epsilon}t^{-2+\epsilon})$, contributes to the large-$m^{2}t$ expansion; if it contributed to the small-$m^{2}t$ expansion, it would exceed the bound for small $t$. To summarize, we should choose $-\epsilon<v_{0}<0$ in this calculation.
@@ C.2 Loop integrals
×$\quad{}\times\bigg{[}\left(\frac{uv+su+sv}{s+v}\right)^{a}\frac{\Gamma(a)\Gamma(d/2-a-b)}{\Gamma(d/2-b)}{}_{2}F_{1}(a;b;1+a+b-d/2;\frac{uv+su+sv}{(s+v)(u+v)})$
@@
×$\qquad{}\qquad{}\times{}_{2}F_{1}(\frac{d}{2}-a,\frac{d}{2}-b;\frac{d}{2}+1-a-b;\frac{uv+su+sv}{(s+v)(u+v)})\bigg{]},$ (87)
@@
×$\quad{}\times\int_{0}^{\infty}\mathrm{d}\beta\beta^{b-1}\frac{1}{(s+v+\beta)^{a}}\frac{1}{[u(v+\beta(1-x))+(s+\beta x)u+(s+\beta x)(v+\beta(1-x))]^{d/2-a}}.$ (88)
@@ Appendix D Perturbative series
$C^{S,\gg 1}_{1,5}$ $=\frac{5C_{\mathrm{F}}}{36864}\bigg{(}746314+1012095L_{m}+480897L_{t}+961794\log{2}-967383\log{3}\bigg{)}.$
@@
Based on the numerical computation explained in sec. 2.5 we provide a grid for dimensionless $\bar{S}$ and $\bar{R}$. The format is $\{\overline{m}^{2}(\mu)t,\frac{t}{\overline{m}(\mu)}\bar{S}\}$ or $\{\overline{m}^{2}(\mu)t,t^{2}\bar{R}\}$, where the perturbative series contain $\alpha_{s}(\mu^{2})$, $\log(t\mu^{2})$, and the color factors as parameters. Also the uncertainties of the perturbative coefficients are labeled by $\delta s[...]$ or $\delta r[...]$. Setting them to 00 (±$\pm 1$) corresponds to the central values (estimated upper/lower bounds) of the perturbative coefficients.

### v2

@@ 2.2 Expressing the VEVs in terms of scalar integrals
[…]
@@ 2.3.1 Laplace transform
where $v_{0}$ is a real number chosen such that the integral of eq. (14) converges for $v=v_{0}$. If $\hat{I}=\mathcal{O}(z^{a_{s}})$ for small $z$ and $I=\mathcal{O}(z^{-a_{l}})$ for large $z$, one may choose $a_{l}<v_{0}<a_{s}$. This also implies that $\tilde{I}(v)$ develops singularities in the $v$-plane which encode the structure of the expansions of $I(z)$ in $z$ and in $1/z$, as we will now explain.
@@ 2.3.3 Two-loop examples
$\quad{}\times\int_{0}^{1}\int_{0}^{1}\mathrm{d}u\,\mathrm{d}u^{\prime}\,u^{\epsilon}(4-u(1-u^{\prime})^{2})^{-\frac{3}{2}+v+\epsilon}{}_{2}F_{1}(1,v-\frac{1}{2};2-\epsilon,\frac{1}{4}u(1-u^{\prime})^{2}),$ (32)
@@
$v$ $=\pm\left(n-\frac{1}{2}\right)\qquad\text{and}\qquad v=n+\frac{3}{2}-\epsilon\,,\qquad n\in\mathbb{N}\,.$ (33)
@@
where $\mathcal{T}_{1,N}g_{a}(y)$ denotes the Taylor expansion of $g_{a}(y)$ around $y=1$ up to power $(1-y)^{N}$, $g_{a}^{(k)}(y)$ is the $k^{\text{th}}$ derivative of $g_{a}(y)$, and $N$ an integer part of v3/2v-3/2 (or $v-3/2\leq N<v-1/2$).
@@
with the integer N=v3/2N=-v-3/2. Note that, since the pre-factor only develops poles at positive $v$, the residues at negative $v$ are given entirely by the Taylor expansion coefficients $h^{(k)}(0)$ of the function $h(y)$ which can be easily evaluated analytically.
@@
This integral involves an additional complication, in that its Laplace transform contains singularities at $v=0$ and $v=-\epsilon$, meaning that $v_{0}$ in eq. (15) has to be chosen carefully. As argued in app. B, the choice $\epsilon<v_{0}<0$ properly separates the small- and large-$(m^{2}t)$ expansions, assuming $\epsilon>0$.
@@ 2.4 Differential equations
$i(t)=\sum_{k=0}^{\infty}\left(\frac{a_{k}}{k+2\epsilon}m^{2k+3}t^{k+2\epsilon}+\frac{b_{k}}{k+\epsilon}m^{2(k-\epsilon)+3}t^{k+\epsilon}+\frac{c_{k}}{k}m^{2(k-2\epsilon)+3}t^{k}\right)+\text{const.}$ (45)
@@ 2.5 Numerical evaluation
Since we expect the asymptotic expansions of the integrals to fail in the intermediate region, we also evaluate the integrals numerically on a one-dimensional grid in $m^{2}t$. This is achieved with the help of the program ftint Harlander:2024vmn . It uses Schwinger parametrization of the momentum integrations in order to map the integrals onto a multi-dimensional hypercube. The integrand becomes the product of polynomials in the Schwinger and flow-time parameters, raised to $d$-dependent powers. Such integrals can be evaluated algorithmically by sector decomposition Binoth:2000ps ; Binoth:2003ak , which expresses them in terms of a Laurent series in d4d-4 whose coefficients are convergent integrals which can be evaluated numerically. For the sector decomposition and the numerical integration, ftint relies on pySecDec Borowka:2015mxa ; Borowka:2017idc ; Borowka:2018goh ; Heinrich:2023til . Evaluation of all integrals to a relative precision of $10^{-8}$ or better on a grid of 200 points takes only a few hours on a currently average computer.
@@ 2.6 Small-flow-time expansion
In addition to a direct calculation of the mass expansion outlined above, one can also determine the small-$m^{2}t$ expansion order-by-order by employing the small-flow-time expansion on the operator level Luscher:2011bx as was briefly mentioned in ref. Lange:2021vqg . This effectively amounts to an operator product expansion with $t\to 0$. For $S(t)$ it reads
@@ 3.1 Perturbative results
The renormalized coupling and mass in the $\overline{\rm MS}$ scheme are given through
@@
$\overline{S}$ $=\langle\bar{\chi}_{\mathrm{R}}(t,x)\chi_{\mathrm{R}}(t,x)\rangle\,$, (57)
@@
Figure 1: The perturbative coefficients of the ratio of $\overline{S}$ (left) and those of $\overline{R}$ (right), obtained using two approaches, i.e., expansions and numerical evaluation, as a function of $\overline{m}^{2}t$, cf. eq. (59). The lines in the region $\overline{m}^{2}t<1$ correspond to the small-$\overline{m}^{2}t$ expansion (blue and purple lines) and those in $\overline{m}^{2}t>1$ to the large-$\overline{m}^{2}t$ expansion (red and orange lines). Since the $\mathcal{O}(\alpha_{\mathrm{s}})$ coefficient also depends on $\mu/\overline{m}$ in addition to $\overline{m}^{2}t$, we show results with a few choices: $\mu/\overline{m}=100$ (purple solid), 55 (purple dashed), 11 (purple dotted), and $\mu/\overline{m}=1$ (orange solid), 22 (orange dashed), $1/2$ (orange dotted).
@@ 3.2 Lattice observables and uncertainty estimates for quark mass extractions
$\alpha_{\mathrm{s}}(M^{2}_{\mathrm{Z}})$ $0,\displaystyle=118$0\,, (70)
[…]
@@
Thus, we can determine $\overline{m}$ only up to $\pm\Delta\overline{m}$, where
@@
Figure 5: Estimate of the expected accuracy on mass determination at the current perturbative order $\mathcal{O}(\alpha_{\mathrm{s}})]$. See the text for details.
In the case of the strange and charm quarks, where the small-$\overline{m}^{2}t$ is more relevant, we see the tendency that $\Delta\overline{m}/\overline{m}$ decreases towards smaller values of $t$, because the perturbative uncertainty decreases in this case. For the strange quark, the effect of the larger perturbative uncertainty for $r_{b}$ compared to $r_{c}$ is partly compensated by the larger derivative, leading to a similar expected accuracy on $\overline{m}$. In the case of the bottom quark, the figure on $r_{b}$ indicates the potential to extract its mass value at 121-2 % already at this perturbative order. We note that these are crude estimates, based on the scale variation around our central scale choices, eq. (68) or (69). Therefore they are understood as order-of-magnitude estimates.
Besides the perturbative uncertainty due to missing higher orders, we also have to take into account the uncertainties from the lattice determinations for a full uncertainty estimate. In addition, there are also non-perturbative corrections which we discuss in the next section. We note, however, that the relevant non-perturbative corrections for $\overline{m}^{2}t\gg 1$ are currently unknown, as mentioned below, and their systematic estimate remains an open problem.
@@ 4 Discussion of non-perturbative effects
where $c_{0,1,2}$ are constants, one can see that the largest non-perturbative contribution, $c_{0}\Lambda_{\rm QCD}^{3}$, which is independent of $m$, disappears due to the derivative in $r_{c}$. Hence, the leading non-perturbative effect reduces to $\mathcal{O}(m\Lambda_{\rm QCD}^{2})$, whose magnitude relative to the perturbative effect $m/t$ reduces to $\mathcal{O}(t\Lambda_{\rm QCD}^{2})$, which is much smaller than one and can be well controlled. As for $1-R/(R|_{m=0})$, the leading non-perturbative effect is $\mathcal{O}(m\Lambda_{\rm QCD}^{3}t^{2})$, which is smaller than that found in $R/(R|_{m=0})$, but still can be large compared with the perturbative effect $m^{2}t$. Based on these observations, we conclude that $r_{c}$ is the most suitable quantity for light quark mass determination among $r_{a,b,c}$ and $1-r_{b}$.
@@ 5 Conclusions
We proposed a method to determine quark masses using the gradient flow by considering ratios of VEVs of flowed quark-bilinear operators. They are scheme and gauge invariant and can be calculated both on the lattice and in perturbation theory. We evaluated these VEVs perturbatively up to two loops, i.e. $\mathcal{O}(\alpha_{\mathrm{s}})$, including finite mass effects. Compared to the already known perturbative results in the small-$\overline{m}^{2}t$ limit, we found that the higher-order terms of the small-$\overline{m}^{2}t$ expansion, the large-$\overline{m}^{2}t$ expansion, and full mass dependence are crucial for studying particularly charm and bottom quarks. The VEV ratios exhibit good perturbative behavior. We also estimated the expected accuracy achieved in the mass determination from the perturbative viewpoint.
@@ Appendix B Choosing $v_{0}$ for the integral $J_{17}$
In the first inequality, we used $1/[(p-k)^{2}+m^{2}]<1/(p-k)^{2}$. The bound (83) is valid as long as $2<d<4$; the $k$-integral is UV divergent for $d\geq 4$ and infrared divergent for $d\leq 2$. Therefore, we assume this range of $d$ and thus $\epsilon>0$. $C$ and $C^{\prime}$ are positive constants. From this bound, we can see that the $v=-\epsilon$ singularity, which gives $\mathcal{O}(m^{-2\epsilon}t^{-2+\epsilon})$, contributes to the large-$m^{2}t$ expansion; if it contributed to the small-$m^{2}t$ expansion, it would exceed the bound for small $t$. To summarize, we should choose $\epsilon<v_{0}<0$ in this calculation.
@@ C.2 Loop integrals
$\quad{}\times\bigg{[}\left(\frac{uv+su+sv}{s+v}\right)^{a}\frac{\Gamma(a)\Gamma(d/2-a-b)}{\Gamma(d/2-b)}{}_{2}F_{1}(a;b;1+a+b-d/2;\frac{uv+su+sv}{(s+v)(u+v)})$
@@
$\qquad{}\qquad{}\times{}_{2}F_{1}(\frac{d}{2}-a,\frac{d}{2}-b;\frac{d}{2}+1-a-b;\frac{uv+su+sv}{(s+v)(u+v)})\bigg{]},$ (87)
@@
$\quad{}\times\int_{0}^{\infty}\mathrm{d}\beta\beta^{b-1}\frac{1}{(s+v+\beta)^{a}}\frac{1}{[u(v+\beta(1-x))+(s+\beta x)u+(s+\beta x)(v+\beta(1-x))]^{d/2-a}}.$ (88)
@@ Appendix D Perturbative series
$C^{S,\gg 1}_{1,5}$ $=\frac{5C_{\mathrm{F}}}{36864}\bigg{(}746314+1012095L_{m}+480897L_{t}+961794\log{2}-967383\log{3}\bigg{)}.$ (96)
@@
Based on the numerical computation explained in sec. 2.5 we provide a grid for dimensionless $\bar{S}$ and $\bar{R}$. The format is $\{\overline{m}^{2}(\mu)t,\frac{t}{\overline{m}(\mu)}\bar{S}\}$ or $\{\overline{m}^{2}(\mu)t,t^{2}\bar{R}\}$, where the perturbative series contain $\alpha_{s}(\mu^{2})$, $\log(t\mu^{2})$, and the color factors as parameters. Also the uncertainties of the perturbative coefficients are labeled by $\delta s[$...] or $\delta r[$...]. Setting them to 0 ($\pm 1$) corresponds to the central values (estimated upper/lower bounds) of the perturbative coefficients.